Bargain Deal, Inc., is a leading retailer specializing in consumer electronics. A condensed income statement and balance sheet for the fiscal year ended January 30, 2016, are shown below. Bargain Deal, Inc. Balance Sheet At January 30, 2016 Assets Current assets Cash and cash equivalents Short-term investnenta Accounts receivable, net Merchandise inventoriea Other current assets 2,136 1,321 1,242 5,067 424 10,190 3,713 Total current assets Long-term assets Total assets $13,903 Liabilities and Shareholders Equity Current İiabilities Accounts payable 5,250 Other current liabilities 9,325 2,248 Total current 1iabilities Long-term liabilities Shareholders equity Total 11abilities and shareholdera equity 13,903 Bargain Deal, Ine. Incone Statenent For the Year Ended January 30, 2016 s in mi11ion) Revenues Costs and expenses 39,608 38,169 1,439 Operating income Incone before income taxes Net income Other income (expense) 1,358 Income tax expense $ 615 Includes $224 of interest expense.
For the Year Ended January 30, 2016 (s in millions) Revenues Costs and expenses $39,608 38,169 1,439 Operating income Other income (expense)* (81) Income before income taxes 1,358 743 Income tax expense Net income $ 615 Includes $224 of interest expense. Required: 1-a. Calculate the current ratio for Bargain Deal for its fiscal year ended January 30, 2016. 1-b. Calculate the acid-test ratio for Bargain Deal for its fiscal year ended January 30, 2016. 1-c. Calculate the debt to equity ratio for Bargain Deal for its fiscal year ended January 30, 2016. 1-d. Calculate the times interest earned ratio for Bargain Deal for its fiscal year ended January 30, 2016. (For all requirements, round your answers to 2 decimal places.) Answer is complete but not entirely correct. Current ratio .09 0.50 4.97 17.77 times a. b Acid-test ratio . Debt to equity ratio 1. Times interest earned d. ratio
Answer #1
Ans. Time Interest Earned Ratio = EBIT / Inetrest Expense
= 1,358+(224-81) / 224
= (1,358+143) / 224
= 1,501/224
= 6.70